# Sourced Carrollian Fluids Dual to Black Hole Horizons

### v2

@@ II.2 The Metric & Gauge Choices: Near Horizon
We focus on the event horizon of a black hole and the dual Carrollian fluid description of it. To work with an event horizon embedded in a 4-dimensional spacetime manifold $\mathscr{M}$, we adopt a version of the general form of the metric in [57]. [^6] Compared to the metric presented in [57], we use a set of coordinates for which the coordinates restricted to the base manifold are the usual coordinates on a sphere, i.e., $z^{A}=\sigma^{A}$, and the Jacobian matrix is given by a Kronecker delta, i.e., $J_{A}^{\hskip 4.0ptB}=\delta_{A}^{B}$. On the other hand, to make contact with the near-horizon metric in Gaussian-null coordinates (as used in [21, 54, 55]), we also set the horizontal shift vector $b_{A}=0$ while allowing for a non-vanishing horizontal velocity field $V^{A}$ [57]. We further rescale the time coordinate to absorb the scale factor $\alpha$ so that it does not explicitly appear in our expressions. [^7] Absorbing $\alpha$ into the time coordinate allowed the system of partial differential equations obtained through the perturbative scheme (section III below) to close. An affine choice of $\alpha$ is further discussed in the Appendix B.1. With these choices, the metric on $\mathscr{M}$ is given by the line element
@@
where $v$ denotes the null time, $r$ is the radial coordinate and $x^{A}$ denotes the coordinates on a 2-sphere, and $V^{2}\equiv q_{AB}V^{A}V^{B}$. Note that $x^{a}=\left(v,r,x^{A}\right)$ are the coordinates on $\mathscr{M}$, $x^{i}=\left(v,x^{A}\right)$ are the coordinates on $\mathscr{H}$ and $\mathscr{N}$, and $x^{A}$ are the coordinates on $\mathcal{S}$, which is the sphere base manifold of $\mathscr{N}$. As we will work with small perturbations of a Schwarzschild black hole in Section V, one can interpret the coordinates as perturbed versions of the advanced Eddington–Finkelstein coordinates.
After adopting the above gauge to obtain the metric, the parameters $\rho$, $V^{A}$, and $q_{AB}$ are still functions of $x^{a}$ in general. Since we are interested in the dynamics of the horizon where $r=0$, we focus on the near-horizon behavior and impose the following expansions:
@@
Thus, the metric we obtain agrees with those of [21, 55] up to $\mathcal{O}\left(r^{2}\right)$. Note here that we impose the limit $r\rightarrow 0$ (and $\rho\rightarrow 0$) to go from the timelike stretched horizons to the null event horizon.
@@ III Horizon Dynamics & Fluid Interpretation
where $N_{i}^{\hskip 4.0ptj}$ is the generalized news tensor and $N$ is its trace [57]. [^8] See [57] and [63] for the definition of the generalized news tensor. Here, we omit a detailed discussion as our goal is to focus on the null hypersurface. On the null surface $\mathscr{N}$, this tensor becomes the Weingarten map of the surface, i.e., $\lim_{\rho\rightarrow 0}N_{i}^{\hskip 4.0ptj}=W_{i}^{\hskip 4.0ptj}$, which is defined by $W_{i}^{\hskip 4.0ptj}\equiv\Pi_{i}^{\hskip 4.0ptc}\left(\nabla_{c}n^{d}\right)\Pi_{d}^{\hskip 4.0ptj}$. This definition is compatible with the limit $\rho\rightarrow 0$ and hence is also valid on the null surface $\mathscr{N}$ [64, 56]. For the gauge we consider in this paper, the generalized news tensor and the Weingarten map also agree on any $\mathscr{H}$.
The projection of the Einstein tensor $G_{ab}$ on $\mathscr{M}$ along $n^{a}$ can be described by the conservation of the stress-energy tensor on $\mathscr{H}$ [56]
@@
On the null surface $\mathscr{N}$, Eq. (26) is still satisfied in vacuum. However, the stress-energy tensor in this case has to be interpreted as a Carrollian stress-energy tensor. The resulting equations are not the equations of relativistic hydrodynamics but those of Carrollian hydrodynamics [32, 56, 57]. Therefore, going from the timelike $\mathscr{H}$ to null $\mathscr{N}$ via the limit $\rho\rightarrow 0$ intuitively corresponds to taking the $c\rightarrow 0$ limit of the relativistic hydrodynamics equations to obtain the Carrollian ones [21].
@@ III.1 Carrollian Fluid/Horizon Dictionary
Parallels between the equations of fluid dynamics and the dynamics of the horizon can be drawn by comparing the Carrollian fluid equations of motion [32, 33] with geometrical (gravitational) equations describing the motion of the null surface in the ambient spacetime. These geometrical equations are commonly referred to as the null Raychaudhuri and Damour–Navier–Stokes (DNS) equations [21, 54]. Below, we express these equations on the horizon in the respective order, imposing the gauge above:
@@
The fluid interpretation of the horizon geometry also follows from the terms appearing in the stress-energy tensor. As constructed using the Weingarten map, the stress-energy tensor defined on $\mathscr{H}$ is given by [57]:
@@
Here, the scalars $\mathcal{E}$ and $\mathcal{P}$ are the energy density and the pressure of the fluid, respectively. The one-form $\pi_{i}$ is interpreted as the fluid momentum density, whereas the vector $\mathcal{J}^{i}$ is the heat flux. The tensor $\tau_{i}^{\hskip 4.0ptj}$ is identified with the viscous stress tensor of the fluid. Evaluating this stress-energy tensor for the black hole horizon [57] The following Carrollian fluid/horizon geometry identifications are made on the horizon:
@@
while $\mathcal{J}^{A}\stackrel{{\scriptstyle\mathscr{N}}}{{=}}0$. The shear viscosity is given by $\eta=1/2$. Note that $\mathcal{P}$ encodes the pressure of the Carrollian fluid and the ambient pressure in the sense that $\mathcal{P}$ is non-zero even when the energy density of the fluid vanishes. It i the surface gravity $\kappa$ that plays the role of the ambient pressure for the Carrollian fluid through this dictionary identification. This definition of pressure agrees with the one in [56]. This dictionary identification need not be modified when the matter fields are introduced in the bulk of the spacetime.
To have a closed system of equations that can be solved consistently, we need two more equations obtained from different projections of the Einstein tensor (e.g. [65, 55]). We choose these components to be the traceless symmetric part of the horizontal projection of the Einstein tensor and the mixed tangential-radial projection, which are expressed on the horizon with our gauge choices as follows:
@@ IV Fluid Sources
Here, the scalar field is defined on $\mathscr{M}$, that is, $\Phi=\Phi(x^{a})$. However, near the horizon, we will assume that $\Phi=\Phi(x^{i})$ as the field should depend on the radial coordinate through the null time $v$ [20]. Plugging this stress-energy tensor in Eq. (32), we obtain the following source term on $\mathscr{N}$:
@@
Explicitly, the existence of the scalar field in the bulk invokes the following contributions to the horizon equations through the field equations:
@@
We observe that for a generic $\Phi\left(v,r,\theta,\phi\right)$, the change of the radial profile of the scalar field, $\partial_{r}\Phi$, plays a role in the last two equations even on the horizon where the stretching vanishes. As expressed above, we will neglect the contributions from $\partial_{r}\Phi$ near the horizon, which will further simplify our treatment.
@@ V Perturbative Treatment
Nonlinear effects have recently garnered renewed attention in black hole perturbation theory, with a focus on analyzing the ringdown phase of gravitational wave signals. Some of these efforts study the imprints of nonlinearities induced by the non-stationary nature of the black hole as it accretes energy and angular momentum (effectively resulting in changes to the black hole parameters) on the ringdown signal [66, 67]. Others make use of second-order black hole perturbation theory to study mode couplings and the generation of quadratic quasinormal modes during ringdown [68, 69, 70, 71, 72, 73]. A recurring theme is the presence of a transient phase whose complexity exceeds that of a simple decomposition into linear modes. This highlights the need for novel approaches to capture the dynamics of this stage and to elucidate the mechanisms governing the transition toward linear behavior on an effectively static background (e.g. [17, 18, 19]).
@@
Black holes interact with their surroundings. The gravitational field of the matter in the bulk affects the dynamics of the horizon. Furthermore, the infalling matter to the black hole results in the expansion and shearing of the horizon. We study such a dynamical scenario by adopting a perturbative scheme. We focus on a Schwarzschild black hole perturbed by an infalling massless scalar field through a small coupling parameter, $\epsilon$. We follow a numerical implementation similar to that in [55], while incorporating matter fields.
@@
The following is the system of equations where the terms are reordered:
  • Order $\mathcal{O}(\epsilon)$:
    $\partial_{v}\Omega^{(1)}_{AB}$ $=2M\partial_{v}\lambda_{AB}^{(1)}+\frac{1}{2}\lambda_{AB}^{(1)}$
    $\quad-\frac{1}{2M}\Omega_{AB}^{(1)}-2M{\nabla}_{\langle A}\mathcal{H}_{B\rangle}^{(1)}$ (54)
    $\kappa^{(1)}$ $=-M{\nabla}_{A}\mathcal{H}^{A(1)}+\frac{M}{2}\hat{R}^{(1)}$ (56)
  • Order $\mathcal{O}(\epsilon^{2})$:
Thus, we have obtained a system of partial differential equations that we solve numerically. Because the quantities we work with are defined on the sphere, we introduce spin-weighted functions and recast the equations in terms of these.
@@ V.1 Spin-weighted Functions
To solve the set of coupled partial differential equations, we follow the spin-weighted formalism [74, 75, 76]. We do so by defining the complex dyad $q^{A}$ which satisfies the following relations:
@@
Note that the perturbations on the induced metric $\Omega_{AB}$ and $\lambda_{AB}$ are assumed to be traceless. Therefore, $\bar{q}^{A}q^{B}\Omega_{AB}=0=\bar{q}^{A}q^{B}\lambda_{AB}$. We also rename the scalar quantities to follow the convention in [55] and for the ease of notation:
@@
Following [75], we compute the curvature scalars on the horizon in terms of the spin-weighted scalars:
@@
Finally, we express all the terms above in terms of spin-weighted spherical harmonics [77, 74]. As an example, we decompose the scalar field as follows:

### v3

@@ II.2 The Metric & Gauge Choices: Near Horizon
We focus on the event horizon of a black hole and the dual Carrollian fluid description of it. To work with an event horizon embedded in a 4-dimensional spacetime manifold $\mathscr{M}$, we adopt a version of the general form of the metric in [freidel2024geometrycarrollianstretchedhorizons]. [^5] Compared to the general metric given by Eq.(62)(62) of [freidel2024geometrycarrollianstretchedhorizons], we use a set of coordinates for which the coordinates restricted to the base manifold are the usual coordinates on a sphere, i.e., $z^{A}=\sigma^{A}$, and the Jacobian matrix is given by a Kronecker delta, i.e., $J_{A}^{\hskip 4.0ptB}=\delta_{A}^{B}$.
On the other hand, to make contact with the near-horizon metric in Gaussian-null coordinates (as used in [Donnay_2019, Adami_2021, Redondo_Yuste_2023]), we also set the Carrollian connection $b_{A}=0$ while allowing for a non-vanishing velocity field $V^{A}$. We further rescale the time coordinate to absorb the scale factor $\alpha$ so that it does not explicitly appear in our expressions. [^6] Absorbing $\alpha$ into the time coordinate allowed the system of partial differential equations obtained through the perturbative scheme (section III below) to close. An affine choice of $\alpha$ is further discussed in the Appendix B.1. With these choices, the metric on $\mathscr{M}$ is given by the line element [^7] Each constant-$r$ slice of this metric can be expressed in the Randers–Papapetrou parametrization [Donnay_2019].
@@
where $v$ denotes the null time, $r$ is the radial coordinate, and $x^{A}$ denotes the coordinates on a 2-sphere, and $V^{2}\equiv q_{AB}V^{A}V^{B}$. Note that $x^{a}=\left(v,r,x^{A}\right)$ are the coordinates on $\mathscr{M}$, $x^{i}=\left(v,x^{A}\right)$ are the coordinates on $\mathscr{H}$ and $\mathscr{N}$, and $x^{A}$ are the coordinates on $\mathcal{S}$, which is the sphere base manifold of $\mathscr{N}$. As we will work with small perturbations of a Schwarzschild black hole in Section V, one can interpret the coordinates as perturbed versions of the advanced Eddington–Finkelstein coordinates.
After adopting the above gauge to obtain the metric, the stretching $\rho$, the horizontal velocity field $V^{A}$, and the sphere metric $q_{AB}$ are still functions of $x^{a}$ in general. Since we are interested in the dynamics of the horizon where $r=0$, we focus on the near-horizon behavior and impose the following expansions [^8] Compare our choices with the Gaussian null coordinates used in Eq.(119-122) of [freidel2024geometrycarrollianstretchedhorizons].:
@@
Here, $\kappa$ is related to the surface gravity, $U^{A}$ is the horizontal vector field, $\Omega_{AB}$ is the sphere metric and $\lambda_{AB}$ are its first-order deformations in the radial coordinate. See Appendix E for explicit examples of these functions in Schwarzschild and slowly-rotating Kerr spacetimes. Thus, the metric we obtain agrees with those of [Donnay_2019, Redondo_Yuste_2023] up to $\mathcal{O}\left(r^{2}\right)$.
$\mathrm{d}s^{2}_{\mathscr{M}}=$ $-2r\kappa\,\mathrm{d}v^{2}+2\mathrm{d}v\,\mathrm{d}r\,-2rU_{A}\,\mathrm{d}v\,\mathrm{d}x^{A}$
$\quad+\left(\Omega_{AB}-r\lambda_{AB}\right)\mathrm{d}x^{A}\,\mathrm{d}x^{B}+\mathcal{O}\left(r^{2}\right)$ (19)
In the stretching formalism, each $r=\mathrm{const.}>0$ represents a timelike hypersurface. Note here that, after imposing the expansions in $r$, the limit $r\rightarrow 0$ (and $\rho\rightarrow 0$) brings the timelike stretched horizons to the null event horizon.
$\mathrm{d}s^{2}_{\mathscr{N}}=0\,\mathrm{d}v^{2}+0\,\mathrm{d}v\,\mathrm{d}x^{A}+\Omega_{AB}\,\mathrm{d}x^{A}\,\mathrm{d}x^{B}$ (20)
We kept above the $v$ direction explicit for pedagogical reasons to emphasize that the event horizon is endowed with a 3-dimensional degenerate (null) metric.
@@ III Horizon Dynamics & Fluid Interpretation
The study of the Carrollian limit began with the work of Lévy-Leblond [Levyleblond1965] and Gupta [Gupta1966] as a mathematical exploration of the contraction [Inonu1953] of the Poincaré group, in which the speed of light vanishes. These studies have evolved into an active field in physics in the last decade [LevyLeblond2022]. Some of the works on the field focused on the relations between the asymptotic symmetries of asymptotically flat spacetimes with Carroll groups and their use in flat holography [Duval_2014_2, Ciambelli_2018_2, Ciambelli_2019, Ciambelli_2019_2, ostendo:fontanella2025carrolllimitadscfttriality]; Carrollian fluids as the limit of relativistic hydrodynamics [Ciambelli_2018, Ciambelli2019, Petkou_2022, Freidel_2023]; the interacting particle motion [Bergshoeff_2014] and dynamics of extended objects [Cardona_2016]; Carrollian limits of relativistic field theories [Basu_2018, Henneaux_2021, Gupta_2021, ostendo:Hao_2022, Hansen_2022, ostendo:Bagchi_2023, ciambelli2023carrollgeodesics, Ecker_2023, Baiguera_2023, ciambelli2023dynamicscarrollianscalarfields] [^9] Note also the early work (1998) by Dautcourt [dautcourt1998ultrarelativisticlimitgeneralrelativity] on the “ultrarelativistic” limit of general relativity and the book (2000) by Klauder [Klauder:2000ud] for ultralocal models in quantum field theories.; applications in cosmology with emphasis on dark energy [deboer2021carrollsymmetrydarkenergy, deboer2023carrollstories].
Thermodynamic properties of Carrollian fluids have also been studied in recent works: for example, in [ostendo:Armas_2024] from the point of view of symmetries (as was developed in a general framework in [ostendo:de_Boer_2018]), and in [ostendo:arenashenriquez2025radiation] within a holographic setup by relating the gravitational radiation with the dissipation in the dual Carrollian theory on the boundary. Despite the progress in the field, Carrollian quantum field theories are not yet fully understood as challenges regarding ill-defined partition functions [deboer2021carrollsymmetrydarkenergy, deboer2023carrollstories], gauge dependence of mass and coupling parameters [Mehra_2023, sharma2025studiescarrollianquantumfield], sensitivity to UV and IR regimes as well as exhibiting a UV/IR mixing [Cotler_2024] remain.
In this section, we discuss the equations that describe the dynamics of the horizon and their interpretation as a hydrodynamical description of a dual Carrollian fluid.
@@
where $N_{i}^{\hskip 4.0ptj}$ is the generalized news tensor and $N$ is its trace [freidel2024geometrycarrollianstretchedhorizons]. [^10] See [freidel2024geometrycarrollianstretchedhorizons] and [Riello_2024] for the definition of the generalized news tensor. Here, we omit a detailed discussion as our goal is to focus on the null hypersurface. On the null surface $\mathscr{N}$, this tensor becomes the Weingarten map of the surface, i.e., $\lim_{\rho\rightarrow 0}N_{i}^{\hskip 4.0ptj}=W_{i}^{\hskip 4.0ptj}$, which is defined by $W_{i}^{\hskip 4.0ptj}\equiv\Pi_{i}^{\hskip 4.0ptc}\left(\nabla_{c}n^{d}\right)\Pi_{d}^{\hskip 4.0ptj}$. This definition is compatible with the limit $\rho\rightarrow 0$ and hence is also valid on the null surface $\mathscr{N}$ [Chandrasekaran_2022, freidel2022carrollian]. For the gauge we consider in this paper, the generalized news tensor and the Weingarten map also agree on any $\mathscr{H}$.
The projection of the Einstein tensor $G_{ab}$ on $\mathscr{M}$ along $n^{a}$ can be described by the conservation of the stress-energy tensor on $\mathscr{H}$ [freidel2022carrollian]
@@
On the null surface $\mathscr{N}$, Eq. (28) is still satisfied in vacuum. However, the stress-energy tensor in this case has to be interpreted as a Carrollian stress-energy tensor. The resulting equations are not the equations of relativistic hydrodynamics but those of Carrollian hydrodynamics [Ciambelli_2018, freidel2022carrollian, freidel2024geometrycarrollianstretchedhorizons]. Therefore, going from the timelike $\mathscr{H}$ to null $\mathscr{N}$ via the limit $\rho\rightarrow 0$ intuitively corresponds to taking the $c\rightarrow 0$ limit of the relativistic hydrodynamics equations to obtain the Carrollian ones [Donnay_2019].
@@ III.1 Carrollian Fluid/Horizon Dictionary
The particular case of Carrolllian fluid dynamics and its dual interpretation as horizon dynamics can be drawn by comparing the Carrollian fluid equations of motion [Ciambelli_2018, Ciambelli2019] with geometrical (gravitational) equations describing the motion of the null surface in the ambient spacetime. These geometrical equations are commonly referred to as the null Raychaudhuri and Damour–Navier–Stokes (DNS) equations [Donnay_2019, Adami_2021]. Below, we express these equations on the horizon in the respective order, imposing the gauge above:
@@
The fluid interpretation of the horizon geometry also follows from the terms appearing in the stress-energy tensor. As constructed using the Weingarten map, the stress-energy tensor defined on $\mathscr{H}$ is given by [freidel2024geometrycarrollianstretchedhorizons]:
@@
Here, the scalars $\mathcal{E}$ and $\mathcal{P}$ are the energy density and the pressure of the fluid, respectively. The one-form $\pi_{i}$ is interpreted as the fluid momentum density, whereas the vector $\mathcal{J}^{i}$ is the heat flux. The tensor $\tau_{i}^{\hskip 4.0ptj}$ is identified with the viscous stress tensor of the fluid. Evaluating this stress-energy tensor for the black hole horizon [freidel2024geometrycarrollianstretchedhorizons] The following Carrollian fluid/horizon geometry identifications are made on the horizon:
@@
while $\mathcal{J}^{A}\stackrel{{\scriptstyle\mathscr{N}}}{{=}}0$. The shear viscosity is given by $\eta=1/2$. Note that $\mathcal{P}$ encodes the pressure of the Carrollian fluid and the ambient pressure in the sense that $\mathcal{P}$ is non-zero even when the energy density of the fluid vanishes. It i the surface gravity $\kappa$ that plays the role of the ambient pressure for the Carrollian fluid through this dictionary identification. This definition of pressure agrees with the one in [freidel2022carrollian]. This dictionary identification need not be modified when the matter fields are introduced in the bulk of the spacetime.
To have a closed system of equations that can be solved consistently, we need two more equations obtained from different projections of the Einstein tensor (e.g. [G_mez_2001, Redondo_Yuste_2023]). We choose these components to be the traceless symmetric part of the horizontal projection of the Einstein tensor and the mixed tangential-radial projection, which are expressed on the horizon with our gauge choices as follows:
@@ IV Fluid Sources
Here, the scalar field is defined on $\mathscr{M}$, that is, $\Phi=\Phi(x^{a})$. However, near the horizon, we will assume that $\Phi=\Phi(x^{i})$ as the field should depend on the radial coordinate through the null time $v$ [Iqbal_2009]. Plugging this stress-energy tensor in Eq. (34), we obtain the following source term on $\mathscr{N}$:
@@
Explicitly, the existence of the scalar field in the bulk invokes the following contributions to the horizon equations through the field equations: [^11] Observe that for a generic $\Phi\left(v,r,\theta,\phi\right)$, the change of the radial profile of the scalar field, $\partial_{r}\Phi$, plays a role in the last two equations even on the horizon where the stretching vanishes. As expressed before, we will neglect the contributions from $\partial_{r}\Phi$ near the horizon, which will further simplify our treatment.
@@ V Perturbative Treatment
Nonlinear effects have recently garnered renewed attention in black hole perturbation theory, with a focus on analyzing the ringdown phase of gravitational wave signals. Some of these efforts study the imprints of nonlinearities induced by the non-stationary nature of the black hole as it accretes energy and angular momentum (effectively resulting in changes to the black hole parameters) on the ringdown signal [Sberna_2022, May_2024]. Others make use of second-order black hole perturbation theory to study mode couplings and the generation of quadratic quasinormal modes during ringdown [Loutrel_2021, Ripley_2021, Ma_2024, Bucciotti_2024, Khera_2023, Khera_2025]. A recurring theme is the presence of a transient phase whose complexity exceeds that of a simple decomposition into linear modes. This highlights the need for novel approaches to capture the dynamics of this stage and to elucidate the mechanisms governing the transition toward linear behavior on an effectively static background (e.g. [Yang_2015, Iuliano:2024ogr, Ma:2025rnv]).
@@
Black holes interact with their surroundings. The gravitational field of the matter in the bulk affects the dynamics of the horizon. Furthermore, the infalling matter to the black hole results in the expansion and shearing of the horizon. We study such a dynamical scenario by adopting a perturbative scheme. We focus on a Schwarzschild black hole perturbed by an infalling massless scalar field through a small coupling parameter, $\epsilon$. We follow a numerical implementation similar to that in [Redondo_Yuste_2023], while incorporating matter fields.
@@
We shall solve the horizon equations (given explicitly in Appendix C) numerically. Because the quantities we work with are defined on the sphere, we introduce spin-weighted functions and recast the equations in terms of these.
@@ V.1 Spin-weighted Functions
To solve the set of coupled partial differential equations, we follow the spin-weighted formalism [Goldberg:1966uu, Gomez_1997, Bishop_1997]. We do so by defining the complex dyad $q^{A}$ which satisfies the following relations:
@@
Note that the perturbations on the induced metric $\Omega_{AB}$ and $\lambda_{AB}$ are assumed to be traceless. Therefore, $\bar{q}^{A}q^{B}\Omega_{AB}=0=\bar{q}^{A}q^{B}\lambda_{AB}$. We also rename the scalar quantities to follow the convention in [Redondo_Yuste_2023] and for the ease of notation:
@@
Following [Gomez_1997], we compute the curvature scalars on the horizon in terms of the spin-weighted scalars:
@@
Finally, we express all the terms above in terms of spin-weighted spherical harmonics [10.1063/1.1931221, Goldberg:1966uu]. As an example, we decompose the scalar field as follows: